What is the total solubility of a weak acid (S) when pH of the solution equals...

  1. What is the total solubility of a weak acid (S) when pH of the solution equals to the pKaof the weak acid. Its So is 0.01M?
Answer #1

Solution:

Suppose weak acid is HA. It dissociated as,

HA === H+ + A-

0.01 M --------0

(0.01 - X) ------ X

According to Hendersen equation,

pH = pka + log [A-] / [HA]

When, pH = pka, then

log [A-] / [HA] = 0

Hence,

[A-] / [HA] =1

Thus, X/ (0.01 -X) = 1

2X = 0.01

X = 0.05 M

When, pH = pka for a weak acid, then acid is half ionized. Thus,

S = 0.01 M /2 = 0.05 M = 0.05 mol/L